OpenStudy (anonymous):

addition with unlike denominators 1/3+1/4=

OpenStudy (amistre64):

make em the same...

OpenStudy (amistre64):

the easiest way is to multiply the bottoms to get a common denominator. might not be the lowest one, but itll work in a pinch

OpenStudy (amistre64):

(4)(3) = 12 1 (4) 1 (3) 4+3 ---- + ---- = ----- 3 (4) 4 (3) 12

OpenStudy (amistre64):

what happens when we multipy nything by (1)? does it change in value?

OpenStudy (amistre64):

the number 4/4 and 3/3 are both equal to (1) and so we can use that to our advantage to get a common denominator. Does that make sense?

OpenStudy (anonymous):

ok i got 12 at the bottom for both but i dont get how to get the two top numbers

OpenStudy (amistre64):

it is a sly trik we use in math. it is to multiply by some form of "1". Tell me is 3/3 = 1? and is 4/4 = 1?

OpenStudy (amistre64):

1 (4) 1 (3) 4+3 -- -- + -- -- = ----- 3 (4) 4 (3) 12 Do you see it?

OpenStudy (anonymous):

but the tow bottom number arent the same in the 2 one u did

OpenStudy (anonymous):

7/12

OpenStudy (anonymous):

amistre 64 the two bottom dumbers arent the same on the second one